OpenStudy (4itc):

what is the product of the complex numbers (-3i+4) and (3i+4) ? A) 1 B) 7 C) 25 D) -7 + 24i E) 7 + 24i

OpenStudy (4itc):

i times-ed the two together and got -9i + 16... maybe i need to add

OpenStudy (4itc):

then id get 8

OpenStudy (mayankdevnani):

use \(\large \bf (a+b)(a-b)=a^2-b^2\) property

OpenStudy (mayankdevnani):

\[\large \bf (4+3i)(4-3i)=4^2-(3i)^2\]

OpenStudy (4itc):

awesome! that helped a lot! at least i know were to start now. thanks!

OpenStudy (4itc):

so far got... 16+12i+12-9i^2 = 16 -9i^2

OpenStudy (mayankdevnani):

correct !

OpenStudy (mayankdevnani):

so what is the value of \(\large \bf i^2\)?

OpenStudy (4itc):

eather (x-1) + (x+1) or add it with the other i^2

OpenStudy (mayankdevnani):

well \[\large \bf i=\sqrt{-1}\]

OpenStudy (mayankdevnani):

so what would be \(\large \bf i^2\) then ?

OpenStudy (4itc):

√-9i^3 = -3 and 3

OpenStudy (4itc):

ops ^2

OpenStudy (mayankdevnani):